OpenStudy (scienceteen8):

1.Carol spends 1/2 hour doing math homework and 4/5 hour doing science homework. About how much time did she spend in all doing her homework? A.about 1 hour B.about 1 1/2 hours C.about 2 hours D.about 2 1/2 hours

OpenStudy (aaronq):

when you add fractions, make the bases (denomintors) the same before adding the numerators Whats a common factor for 2 and 5?

OpenStudy (scienceteen8):

1?

OpenStudy (aaronq):

sure, but you cant multiply them by 1 to make them the same number, right? 1 times 2 is still just 2. the point is to make the bases the same

OpenStudy (aaronq):

The easiest way (and will always work) is just to multiply a fraction by the denominator of the other. In this case we have \(\sf \dfrac{1}{2}+\dfrac{4}{5}\), so, \(\sf 5*\dfrac{1}{2}+2*\dfrac{4}{5}\) \( \sf \dfrac{5}{10}+\dfrac{8}{10}\) we can just add these numerators togehter now and simply (from an improper to a proper fraction) to get one of the answers

OpenStudy (aaronq):

lol